OpenStudy (anonymous):

What 2 numbers multiply into 2,475

OpenStudy (mayankdevnani):

1 number =45*55=2475

OpenStudy (anonymous):

thnx people :) got the answer

OpenStudy (mayankdevnani):

2 number=2475*1=2475

OpenStudy (mayankdevnani):

welcome

OpenStudy (anonymous):

what about 495?

OpenStudy (mayankdevnani):

1 no.=495*1=495

OpenStudy (mayankdevnani):

2 number= 5*98=495

OpenStudy (mayankdevnani):

@SchoolFreak

OpenStudy (anonymous):

two numbers that go into 99?

OpenStudy (mayankdevnani):

1 number=1*99=99 2 number=3*33=99

OpenStudy (mayankdevnani):

@SchoolFreak

OpenStudy (anonymous):

kk thnx

OpenStudy (mayankdevnani):

welcome